OpenStudy (anonymous):

in slope intersept form wha is the anser to 3x+5y=-5

OpenStudy (anonymous):

y=-3/5+1 (subtract 3x from both sides, divide both sides by 5 to get y on its own)

OpenStudy (anonymous):

sorry, it's -1, not +1.

OpenStudy (lieutenantgeneral):

5y = 3x + 5 y = (3/5)x + 1 The slope is 3/5 and the y-intercept is 1.